Grade 9General Physics

A body of mass 10kg droped from a height of 10m collide with earth surface . find the velocity of the body after colliding

Adarsh kumar
10 Years agoGrade 9

1 Answer

bharath
10 Years ago
m1v1=mv2
as initial velocity is zero there exists final velocity.
mgh1=mgh2
mgh1+1/2mv1^2=Mgh+1/2mv2^2
gh1+1/2mv1^2=gh2+1/2mv2^2
gh1=1/2mv2^2
9.8*10=5*v2^2
v2^2=98/5
v2=4.27